# from Monocular RGB-D Video

### v1

@@ 4.3 Analysis
We further analyze different components of our proposed method by conducting a set of ablation experiments on the 9 real-world scenes with quantitative results shown in Tab. 3 and qualitative demonstrations in Fig. 4-7.
Temporal View-dependent SDS. We start by analyzing the diffusion prior. As in Tab. 3, our full model (last row) exhibits a significant improvement in CLIP similarity in 360° compared to the variant without diffusion priors (first row). Notably, the results also highlight the diffusion priors as a regularization for surface reconstruction in the observed region, leading to the improvement in reconstruction accuracy (0.88 v.s. 0.99) by effectively eliminating spurious surfaces.
We further analyze a variant without the temporal condition (second row), where the first frame is used as the condition for the diffusion priors all the time. Results show a degradation in both 3D metrics and CLIP similarity, demonstrating the importance of accurate temporal conditions for estimating the scene geometry when the target object is in motion over time. However, as opposed to using a given frame as a direct data-term supervision, the use of SDS with slightly inconsistent view conditions could still achieve realistic scene completion without drastically affecting real-view reconstruction results.
@@
Effect of Depth Measurement. As in Tab. 3, despite we could achieve competitive CLIP similarity without depth measurement, it fails to capture accurate motion and geometry from the monocular RGB video. We further illustrate the pitfalls of the RGB-based methods in Fig. 6. Magic123 [49] is trained on a single frame with a monocular depth prior as a reference. Due to the scale ambiguity, it is very challenging to recover the accurate geometry (e.g., the human arm) for RGB-based methods.
@@
Effect of Representations with a Canonical Shape. To evaluate the impact of the canonical space, we compare our full model to a canonical space-free representation (w/o canonical space). In the latter, the world space is directly modeled as a 4D space-time grid, which is decomposed into a quad-cube representation with 4 3D hash grids. Without the implicit regularization provided by a canonical space, the quad-cube demonstrates worse temporal consistency in both geometry and appearance reconstruction, as reflected in the quantitative results in Tab. 3.
@@
Limitations. While we demonstrate photo-realistic 360° reconstruction, certain limitations are inherent in our current formulation. Specifically, real-world captures often present challenges such as incomplete views, motion blur, and intricate articulated poses of the target object (see Fig. 8). Our model, which relies on the diffusion prior conditioned on pure RGB frames, may fail in these scenarios and result in incoherent reconstruction, missing geometry, etc. In addition, target objects might also have complex motion in the self-occluded region, which is difficult to model due to the lack of motion priors. Introducing better diffusion priors (e.g., diffusion models that are conditioned on RGB-D images, time, etc) and motion priors can be a promising avenue for future research.

### v2

@@ 4.3 Analysis
We further analyze different components of our proposed method by conducting a set of ablation experiments on the 9 real-world scenes with quantitative results shown in Tab. 3 and qualitative demonstrations in Fig. 5-8.
Temporal View-dependent SDS. We start by analyzing the diffusion prior. As in Tab. 3, our full model (last row) exhibits a significant improvement in CLIP similarity compared to the variant without diffusion priors (first row). Notably, the results also highlight the diffusion priors as a regularization for surface reconstruction in the observed region, leading to the improvement in reconstruction accuracy (0.88 v.s. 0.99) by effectively eliminating spurious surfaces.
We further analyze a variant without the temporal condition (second row), where only the first frame is used as the condition for diffusion prior. Results show a degradation in both 3D metrics and CLIP similarity, demonstrating the importance of accurate temporal conditions for estimating the scene geometry when target object is in motion over time. However, as opposed to using a given frame as a direct data-term supervision, the use of SDS with slightly inconsistent view conditions could still achieve realistic scene completion without drastically affecting real-view reconstruction.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Train view GT view NDR Ours
Figure 4: Qualitative results of novel view synthesis. Please refer to suppl. material for more quantitative & qualitative comparisons.
@@
Effect of Depth Measurement. As in Tab. 3, although our model achieves competitive CLIP similarity without depth measurement, it fails to capture accurate motion and geometry. We further illustrate the pitfalls of RGB-based methods in Fig. 7. Magic123 [57] is trained on a single frame with a monocular depth prior as a reference. Due to the scale ambiguity, it is very challenging to recover the accurate geometry (e.g., the human arm) for RGB-based methods.
@@
Effect of Representations with a Canonical Shape. To evaluate the impact of the canonical space, we compare our full model to a canonical space-free representation (w/o canonical space). In the latter, the world space is directly modeled as a 4D space-time grid, which is decomposed into a quad-cube representation with 4 3D hash grids. Without the implicit regularization provided by a canonical space, the quad-cube demonstrates worse temporal consistency in both geometry and appearance reconstruction as in Tab. 3.
@@
Limitations. While we demonstrate photo-realistic 360° reconstruction, certain limitations are inherent in our current formulation. Specifically, real-world captures often present challenges such as incomplete views, motion blur, and intricate articulated poses of the target object (see Fig. 9). Our model, which relies on the diffusion prior conditioned on pure RGB frames, may fail in these scenarios and result in incoherent reconstruction or missing geometry. Moreover, our model cannot reconstruct complex motion in self-occluded regions due to the absence of motion priors Introducing better diffusion priors (e.g., diffusion models that are conditioned on RGB-D images, time, etc) and motion priors can be a promising avenue for future research.